# GCLS2: Towards Efficient Community Detection using Graph Contrastive Learning with Structure Semantics

### v1

@@ V-D The Hyperparameter Analysis – RQ3
V-E The Case Study
We evaluated the community detection performance of the GCLS2 model on Cora by the T-SNE [35] downscaling. Specifically, after the training of Algorithm 1 for $\varepsilon$ = {10, 30, 500}, respectively, we visualized the distribution of node features $Z$ in two dimensions, using nodes in 7 different colors to represent the 7 communities. The results are shown in Figure 5. We can clearly see that from $\varepsilon$ = 10 to 30 iterations, the labeled nodes are closing in on the communities corresponding to the labels. By 500 iterations, the detection accuracy is significantly improved, validating the effectiveness of our GCLS2 model.
@@ VI Conclusion
In this paper, we analyze the limitations of traditional GCL methods for community detection, and propose a novel Graph Contrastive Learning with Structure Semantics (CGLS2) framework for the community detection. We construct a high-level structure view based on classical community structure, and then extract structural and attribute semantics through the structure similarity semantics encoder to obtain a comprehensive node feature representation. We also design the structure contrastive learning to enhance the structural feature representation of nodes for more accurate community detection. Extensive experiments confirm the effectiveness of our GCLS2 approach.

### v2

@@ IV-D The HGP Algorithm Evaluation – RQ3
IV-E The Hyperparameter Analysis – RQ4
Figure 7 shows the sensitivity analysis of the hyperparameters (structure semantic dimension $d$ and temperature parameter $\tau$) of GCLS2 on Cora and Email-Eu datasets. Then we can answer RQ4. From the results, the ACC, NMI, and MF1 metrics are stable on Cora and Email-Eu datasets for $d$={8, 16, 32, 64, 128}. However, the temperature parameter $\tau$ has the best performance when the $\tau$=1. Since the $\tau$ affects the vector distance of the sample pairs for structure contrastive learning, a larger $\tau$ reduces the effect of structure contrastive learning, resulting in the reduction of ACC, NMI, and MF1 metrics for both datasets.
IV-F The Case Study
We evaluate the community detection performance of the GCLS2 model on Cora by the T-SNE [48] downscaling. Specifically, after the training of Algorithm 2 for $\varepsilon$ = {10, 30, 500}, respectively, we visualize the distribution of node features $Z$ in two dimensions, using nodes in 7 different colors to represent the 7 communities. The results are shown in Figure 8. We can clearly see that from $\varepsilon$ = 10 to 30 iterations, the labeled nodes are closing in on the communities corresponding to the labels. By 500 iterations, the detection accuracy is significantly improved, validating the effectiveness of our GCLS2 model.
V Related Work
Graph Contrastive Learning: The graph contrastive learning optimizes the representation of nodes, by maximizing the agreement between pairs of positive samples and minimizing the agreement between pairs of negative samples to be adapted to various downstream tasks. For the node level, there are two general contrastive ways: i) node-graph, for example, DGI [43] contrasts node feature embeddings of augmented view and original data graph with graph feature embeddings of the original data graph, and MVGRL [44] contrasts node feature embeddings of one view with the graph feature embeddings of another view, ii) node-node, for example, GRACE [11] contrasts node feature embeddings of two augmented views with each other, and NCLA [34] contrasts neighbors’ feature embeddings of multiple augmented views. Although these contrastive learning methods have made greater progress in the feature representation, most contrasted samples in the community detection task are against the community’s inherent information representation, where feature embedding representations within communities should be similar and feature embedding representations between communities should be dissimilar. Our proposed GCLS2 approach uses a high-level structure adjacency matrix as a signal to guide the anchor closer to dense intra-community and away from the inter-community and achieves good detection results even when using community structure information only.
To the best of our knowledge, no prior works proposed to use the structure contrastive loss for the community detection task. The ability to operate without relying on real labels and attributes makes our GCLS2 model highly versatile for real-world applications. Moreover, we present our theoretical analysis for structural contrast loss training. We also design a graph partitioning algorithm to scale our proposed approach to handle large-scale graphs and training can be accelerated by parallel processing.
@@ VI Conclusion
In this paper, we analyze the limitations of traditional GCL methods for community detection, and propose a novel Graph Contrastive Learning with Structure Semantics (CGLS2) framework for the community detection. We construct a high-level structure view based on classical community structure, and then extract structural and attribute semantics through the structure similarity semantics encoder to obtain a comprehensive node feature representation. We also design the structure contrastive learning to enhance the structural feature representation of nodes for more accurate community detection. Moreover, we propose the HGP algorithm, which makes it possible to perform online training and community detection over large-scale graph datasets on low-memory machines. Extensive experiments confirm the effectiveness of our GCLS2 approach and HGP algorithm for the community detection and large-scale graph partitioning.